# Antipodal self-duality of square fishnet graphs

### v1

@@ Abstract
In strongly deformed planar $\mathcal{N}=4$ super-Yang-Mills theory, or fishnet theory, a correlation function of four dimension-$m$ protected operators is given by a single Feynman integral, which involves integrating over locations of a $m\times m$ grid of points. We show that for any integer $m$ this square fishnet graph is invariant under the combined action of a kinematic map and the antipode map of the Hopf algebra on multiple polylogarithms, i.e. it possesses an antipodal self-duality.
@@
where $N_{c}$ is the number of colors and $g^{2}$ is the (’t Hooft) coupling constant. The four-point correlator $\langle X(x_{1})^{n}Z(x_{3})^{m}\bar{X}(x_{2})^{n}\bar{Z}(x_{4})^{m}\rangle$ is given by a single Feynman integral $G_{m,n}$, called a fishnet integral (see fig. 1). We will show that square fishnet integrals with $n=m$ are antipodally self-dual. Unlike previous observations of antipodal dualities, we can provide a proof for all “loop orders”, i.e. all values of $m$. Moreover, the antipodal symmetry can be realized fully at the function level, and we show how the kinematic map on symbol letters arises in this case.
Figure 1: The rectangular fishnet graph for the four-point correlation function $\langle X^{n}\,Z^{m}\bar{X}^{n}\,\bar{Z}^{m}\rangle$, or $G_{m,n}$. The pink lines provide a dual, scattering interpretation.
Our letter is organized as follows: We first provide some background on fishnet integrals and the antipode. We then define the antipodal symmetry which leaves square fishnet integrals invariant and we present a complete proof of the antipodal self-duality. Finally we draw our conclusions. We include some supplemental material where we collect technical details of the proof omitted in the main text.
@@ III Proof of antipodal self-duality
We can easily check that (14) holds for $p=1$, so we only need to prove it for $p>1$. A complete proof of (14) is given in the supplementary material. Here we sketch a simpler argument why (14) holds, although this argument only allows us to show that (14) holds modulo terms proportional to $(2\pi i)^{2}$.
@@
This translates into the fact that the symbol of $\operatorname{Disc}_{1}\operatorname{{\bf\hat{S}}}(f_{p})$ vanishes up to terms proportional to $(2\pi i)^{2}$, in agreement with (17). For the discontinuity at 0, we employ the following identity, whose proof is presented in the supplemental material,
@@ IV Discussion
Many conformally-invariant four-point functions can be evaluated in terms of the same class of functions as the fishnet integrals, namely single-valued (harmonic) polylogarithms. An important class is large $R$-charge correlators, where other determinants of ladder integrals appear [43]. Our results immediately beg the question of whether other classes of correlators or Feynman integrals enjoy an antipodal symmetry. We have already mentioned that fishnet integrals $\phi_{m,n}$ with $n\neq m$ are not antipodally self-dual, which can be established independently of the kinematic map for small $m,n$ by counting dimensions of spaces of coproducts. Also, the generalized fishnet integrals of [43] that we inspected do not possess antipodal self-duality for the specific map in $\operatorname{{\bf\hat{S}}}$.
We also asked whether other polynomials in the ladder integrals $f_{p}$ could be antipodally self-dual under $\operatorname{{\bf\hat{S}}}$, and also satisfy the Steinmann relations [44, 45, 46], which imply that the first two entries in the symbol (or their complex conjugates) cannot both be $1-z$. (The square and rectangular fishnets and the integrals of [43] all satisfy the Steinmann relations.) We found that these conditions are very restrictive, and there are very few such polynomials at small weights. In particular, there is a unique infinite family of polynomials of degree 2 of this type, given by
@@ A Proof of (19)
In the third step we used $(j+1)-2(j-p+1)=2p-j-1$.
@@ B Proof of (14)
In the last step we replaced the upper summation limit $\min(l,p)$ by $l$. This is justified in our case, because the summand contains a factor $\tfrac{1}{(l-k)!}$, which vanishes for $k>l$. We then find

### v2

@@ Abstract
In strongly-deformed planar $\mathcal{N}=4$ super-Yang-Mills theory, or fishnet theory, a point-split single-trace correlation function of four dimension-$m$ scalar operators is given by a single Feynman integral, which involves integrating over locations of a $m\times m$ grid of points. We show that for any integer $m$ this square fishnet graph is invariant under the combined action of a kinematic map and the antipode map of the Hopf algebra on multiple polylogarithms, i.e. it possesses an antipodal self-duality.
@@
where $N_{c}$ is the number of colors and $g^{2}$ is the (’t Hooft) coupling constant. The four-point correlator $\langle\operatorname{Tr}[X^{n}(x_{1})Z^{m}(x_{3})\bar{X}^{n}(x_{2})\bar{Z}^{m}(x_{4})]\rangle$ is given by a single Feynman integral $G_{m,n}$, called a fishnet integral (see fig. 1). We will show that square fishnet integrals with $n=m$ are antipodally self-dual. Unlike previous observations of antipodal dualities, we can provide a proof for all “loop orders”, i.e. all values of $m$. Moreover, the antipodal symmetry can be realized fully at the function level, and we show how the kinematic map on symbol letters arises in this case.
Figure 1: The rectangular fishnet graph for the four-point correlation function $\langle\operatorname{Tr}[X^{n}\,Z^{m}\bar{X}^{n}\,\bar{Z}^{m}]\rangle$, or $G_{m,n}$. The pink lines provide a dual, scattering interpretation.
Our letter is organized as follows: We first provide some background on fishnet integrals and the antipode. We then define the antipodal symmetry which leaves square fishnet integrals invariant and we present a complete proof of the antipodal self-duality. Finally we draw our conclusions. We include two appendices where we collect technical details of the proof omitted in the main text.
@@ III Proof of antipodal self-duality
We can easily check that (14) holds for $p=1$, so we only need to prove it for $p>1$. A complete proof of (14) is given in Appendix B. Here we sketch a simpler argument why (14) holds, although this argument only allows us to show that (14) holds modulo terms proportional to $(2\pi i)^{2}$.
@@
This translates into the fact that the symbol of $\operatorname{Disc}_{1}\operatorname{{\bf\hat{S}}}(f_{p})$ vanishes up to terms proportional to $(2\pi i)^{2}$, in agreement with (17). For the discontinuity at 0, we employ the following identity, whose proof is presented in Appendix A (see also [38, 39]),
@@ IV Discussion
Many conformally-invariant four-point functions can be evaluated in terms of the same class of functions as the fishnet integrals, namely single-valued (harmonic) polylogarithms. An important class is large $R$-charge correlators, where other determinants of ladder integrals appear [45, 46]. Our results immediately beg the question of whether other classes of correlators or Feynman integrals enjoy an antipodal symmetry. We have already mentioned that fishnet integrals $\phi_{m,n}$ with $n\neq m$ are not antipodally self-dual, which can be established independently of the kinematic map for small $m,n$ by counting dimensions of spaces of coproducts. Also, the generalized fishnet integrals of [45, 46] that we inspected do not possess antipodal self-duality for the specific map in $\operatorname{{\bf\hat{S}}}$.
We also asked whether other polynomials in the ladder integrals $f_{p}$ could be antipodally self-dual under $\operatorname{{\bf\hat{S}}}$, and also satisfy the Steinmann relations [47, 48, 49], which imply that the first two entries in the symbol (or their complex conjugates) cannot both be $1-z$. (The square and rectangular fishnets and the integrals of [45, 46] all satisfy the Steinmann relations.) We found that these conditions are very restrictive, and there are very few such polynomials at small weights. In particular, there is a unique infinite family of polynomials of degree 2 of this type, given by
@@ A Proof of (19)
In the third step we used $(j+1)-2(j-p+1)=2p-j-1$. Eq. (19) has also appeared in the context of the expectation value of the charge operator for a free massive boson at finite temperature with a chemical potential [38, 39].
@@ B Proof of (14)
In the last step we replaced the upper summation limit $\min(l,p)$ by $l$. This is justified in our case, because the summand contains a factor $\tfrac{1}{(p-k)!}$, which vanishes for $k>p$. We then find